OpenStudy (anonymous):

Find the limit as x approaches pi/2 for ((sinx/cos(^2)x)-tan(^2)x)

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to \pi/2}\frac{\sin x}{\cos^2x}-\tan^2 x\] Mmmm thinking.

OpenStudy (aum):

L'Hospital's rule is okay to use?

OpenStudy (aum):

\[ \large \lim_{x\to \pi/2}\frac{\sin x}{\cos^2x}-\tan^2 x = \large \lim_{x\to \pi/2}\frac{\sin x}{\cos^2x}-\frac{\sin^2 x}{\cos^2 x} = \\ \large \lim_{x\to \pi/2}\frac{\sin x(1-\sin x)}{\cos^2x} = ? \]

OpenStudy (aum):

0/0 form. Try L'H a couple of times.

OpenStudy (anonymous):

unfortunately we cannot use l'hospital rule

OpenStudy (anonymous):

is there any other way to solve it? this one has me stumped!

geerky42 (geerky42):

To finish where @aum stopped... \[\lim_{x\rightarrow\pi/2}\dfrac{\sin x(1-\sin x)}{\cos^2x}\\~\\\lim_{x\rightarrow\pi/2}\dfrac{\sin x(1-\sin x)}{1-\sin^2x}\\~\\\lim_{x\rightarrow\pi/2}\dfrac{\sin x(1-\sin x)}{(1+\sin x)(1-\sin x)}\\~\\\lim_{x\rightarrow\pi/2}\dfrac{\sin x}{1+\sin x}\]

geerky42 (geerky42):

From here, you can make direct substitution.

OpenStudy (anonymous):

Is geerky42's solution using L'hopital's rule, or is it without?

geerky42 (geerky42):

Without.

OpenStudy (anonymous):

Without. Its just clever factoring.

OpenStudy (anonymous):

ok thank you! for your help!

geerky42 (geerky42):

I replaced \(\cos^2x \) to \(1-\sin^2 x\) because \(\sin^2 x + \cos^2 x = 1\)

geerky42 (geerky42):

Welcome

OpenStudy (anonymous):

what did you replace for tan^2x

geerky42 (geerky42):

\(\dfrac{\sin^2 x}{\cos^2 x}\)

OpenStudy (anonymous):

What happened to the sin^2x from the tan^2x?

geerky42 (geerky42):

\[\lim_{x\rightarrow\pi/2}\dfrac{\sin x}{\cos^2x} - \dfrac{\sin^2x}{\cos^2x} = \lim_{x\rightarrow\pi/2}\dfrac{\sin x-\sin^2 x}{\cos^2x}\] Then we factored out \(\sin x\) in numerator to get this: \[\lim_{x\rightarrow\pi/2}\dfrac{\sin x(1-\sin x)}{\cos^2x}\] Does that help?

OpenStudy (anonymous):

Yes it does thank you, again!

geerky42 (geerky42):

No problem